OpenStudy (anonymous):

Exponential Functions: Can someone show me how to solve this, with or without a calculator: 100 (1-e^(-0.08 t))-70

OpenStudy (anonymous):

\[0=100(1-e ^{-0.08t})-70\] a little cleaner

OpenStudy (rob1525):

\[70=100 - 100e^-0.08t\]

OpenStudy (rob1525):

subtract 100 and divide by -100

OpenStudy (anonymous):

the question is asking to solve for zero to get a decimal

OpenStudy (rob1525):

you can find it with a graph or algebraicly

OpenStudy (anonymous):

i have the answer but i dont understand how they go it 15.04966

OpenStudy (anonymous):

graph it by pluggin in numbers for T?

OpenStudy (rob1525):

Graph it with the t and then find when it crosses the x axis

OpenStudy (anonymous):

oh yes i see what you mea, that point will give me the answer

OpenStudy (rob1525):

It sure will, would you like me to solve algebraicly?

OpenStudy (anonymous):

yes please, i wont have time to graph if I take a test

OpenStudy (rob1525):

\[70-100=-100*e^-0.08t\]

OpenStudy (rob1525):

\[-30/-100=e^-0.08t\]

OpenStudy (anonymous):

there is more?

OpenStudy (rob1525):

sorry got disconnected. Yes.

OpenStudy (rob1525):

Now use the natural log to bring down the exponent. the natural log undoes e.\[\ln(3/10)=\ln(e^-0.08t)\]

OpenStudy (rob1525):

t= ln(3/10)/-0.08 = 15.049

OpenStudy (anonymous):

ok go it, I'm gonna right this down just to make sure i got it thank you!

OpenStudy (rob1525):

no prob ;)